# BuildingView: Constructing Urban Building Exteriors Databases with Street View Imagery and Multimodal Large Language Model

### v1

@@ Abstract.
Building exteriors, Street-View imagery, Multimodal large language model, GeoAI, Urban analytics
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
@@
Urban environments are increasingly becoming the focus of interdisciplinary research, where the integration of spatial and visual data is essential for understanding various aspects of urban life. Within this context, the study of building exteriors has gained significant attention due to its direct impact on critical areas such as energy efficiency(parking-lots, ), environmental sustainability(low-carbon, ), and human-centric urban design(human-perception, ). Building exteriors serve not only as the physical interface between a structure and its environment but also play a crucial role in determining the overall energy performance and aesthetic appeal of urban spaces. Street view imagery, in particular, has emerged as a powerful tool in this domain, offering a unique and detailed perspective on building exteriors that traditional remote sensing methods often overlook(liang2024evaluating, ). Unlike remote sensing imagery, which provides a top-down view of urban landscapes, street view imagery captures buildings at eye level, revealing intricate details of their spatial attributes, materials, and architectural styles(biljecki2021street, ). This granular level of detail is vital for assessing the impact of building exteriors on factors such as urban heat islands, energy consumption, and the overall aesthetic coherence of neighborhoods. While street view imagery offers these advantages, its potential remains underutilized in urban research, particularly when it comes to integrating this visual data with other spatial datasets to provide a comprehensive understanding of building exteriors and their broader urban implications. This highlights the growing need to harness street view imagery more effectively to address the multifaceted challenges of modern urban environments.
@@
Through a systematic search of the urban building exteriors research, it is evident that there is a significant gap in the availability of comprehensive databases that integrate street view imagery with detailed building exterior information. Existing studies tend to focus on either the spatial characteristics of buildings—such as their geographic positioning and interactions with surrounding infrastructure—or the aesthetic and material aspects, including innovations in architectural design and material usage(building-style, )(material, ). However, these perspectives are rarely combined to offer a full understanding of urban building exteriors. The lack of integrated databases that unify spatial data with visual and material information limits the ability to conduct in-depth analyses necessary for informed decision-making in urban planning, architectural design, and environmental policy. Additionally, while recent advancements have leveraged street view imagery for various urban analyses, they primarily focus on environmental factors or general urban features rather than systematically extracting and organizing building exterior information(rundle2011using, ). The integration of large language models with street view imagery for detailed urban analysis remains in its nascent stages, highlighting the need for a more integrated approach to understanding urban building exteriors.
@@
To address these knowledge deficits, we propose BuildingView, an innovative approach that integrates Street View imagery with multimodal large language models to create a comprehensive urban building exteriors database. By leveraging high-resolution visual data from Google Street View, combined with spatial and structural data obtained through the Overpass API and OpenStreetMap, this research sets a new standard for the integration of visual and spatial data in urban informatics, significantly advancing the field. Our key contributions include: (1) the systematic collection and organization of urban building exteriors indicators, providing a comprehensive understanding of critical factors such as energy efficiency, environmental sustainability, and human-centric design. (2) the development of BuildingView, a reusable tool that utilizes Street View imagery and multimodal large language models to construct detailed urban building exterior databases for any urban area globally. (3) the construction of extensive urban building exterior databases for New York, Amsterdam and Singapore, enabling more informed decision-making in urban planning, architectural design, and environmental policy.
2. Related Work
2.1. Indicators of Building Exteriors Indicators
@@
In the realm of urban informatics and geoscience, the development of spatiotemporal databases pertaining to buildings has garnered significant attention. Researchers within the disciplines of urban planning and geography typically concentrate on the spatial characteristics of buildings, such as their geographic positioning and the dynamics of their interaction with adjacent infrastructure. For instance, the architectural and urban planning processes must meticulously consider the spatial interplay between buildings and their adjacent parking facilities to optimize energy efficiency (parking-lots, ).In contrast, architects and designers are more inclined to examine the intrinsic elements of buildings, including materials and aesthetic design. Recently, the architectural style and the materials used for the exterior of buildings are continually subject to innovation and enhancement (building-style, ; material, ).
@@
The academic discourse has increasingly pivoted towards the examination of the connections between buildings and other indicators, with a particular emphasis on their impact on energy consumption and environmental sustainability. Especially, researchers are actively exploring ways to harness more energy from natural sources through building design, thereby reducing the energy costs associated with buildings. Additionally, they are working to mitigate the thermal impacts caused by the urban heat island effect, which is exacerbated by the heat generated within urban environments (energy, ). Concurrently, there is a burgeoning interest in the human-centric aspects of building design, particularly how it fosters engagement and interaction with the public from a participatory perspective (human-perception, ).
@@
While numerous open-source databases exist that cater to specific domains of building information, they are often limited to a single aspect of building data. The foundational building attribute knowledge within these databases is typically derived from field surveys or footprint data extracted from remote sensing imagery (chen2022large, ). To date, no open-source database has been developed that integrates street view imagery to provide a comprehensive, visually-driven repository encompassing the spatial. Thus, we want to build a spatial database with fundamental, and external surface properties of buildings, with a focus on their relevance to energy, environmental, and human-centric dimensions.
2.2. SVI and LLMs in Urban Building Analysis
@@
The advent of street view imagery (SVI) has revolutionized the way urban environments are analyzed, offering a rich and dynamic geospatial data source that rivals traditional remote sensing methods such as satellite imagery (biljecki2021street, ). The comprehensive visual data captured by SVI allows for detailed analysis and understanding of urban environments at the street level, providing critical insights into various urban characteristics and phenomena.
@@
Previous research has demonstrated the utility of SVI in urban studies by integrating it with other urban data sources such as street networks, building information, demographic and socioeconomic data, and survey responses. For instance, SVI has been utilized to audit neighborhood environments (rundle2011using, ), and analyze the effects of green view indices on pedestrian activity(ki2021analyzing, ). Additionally, studies have explored the use of SVI and machine learning to measure the built environment’s impact on crime (hipp2021measuring, ), as well as investigating the relationship between neighborhood walkability factors and walking behaviors using big data approaches (koo2022neighborhood, ). Furthermore, SVI has been applied at a global scale, as seen in the creation of extensive datasets that cover millions of street-level images across hundreds of cities, facilitating large-scale urban science and analytics(hou2024global, ).
@@ 3. Method
In this study, we introduce BuildingView, a novel framework for constructing urban building exteriors databases using Street View imagery and multimodal large language models. The research framework is divided into three primary steps, as illustrated in Figure 1: (1) a systematic literature review of building exterior indicators; (2) sampling of buildings and street view imagery using the Overpass and Google Street View APIs to gather data points; and (3) annotation of street view images using the ChatGPT-4O API to create the Urban Building Exteriors Database. Further details on each of these steps are provided below.
Figure 1. The Workflow of BuildingView
[…]
A comprehensive workflow for analyzing building exteriors using a systematic literature review, sampling of buildings and street view imagery, and annotation with a multimodal large language model, integrating multiple APIs and exporting data in various spatial formats.
3.1. Systematic Literature Review of Building Exteriors Indicators
@@
Global building morphology indicators database has concluded many indicators for urban data science (global-indicators, ). However, a database of buildings directly covering energy, green, and human-centred buildings has not yet been created. Our goal is to establish a building database that encompasses the majority of indicators of interest to researchers. This database will include comprehensive spatial information on buildings and knowledge extracted from street view images through advanced vision analytics and vision-based reasoning techniques.
@@
For that, we have done a systematic literature review on building exteriors Figure 2. The initial step is to identify emerging topics concerning building exteriors over the past five years. To achieve this, we conduct a keyword search on ”building envelope,” ”exterior of building,” and ”building exterior” within the Web of Science database. Our search yields 6,28562856,2856 , 285 relevant scholarly works. There is a growing discussion about the intelligent design of external surfaces of buildings. The integration of innovative materials and intelligent technologies has the potential to significantly enhance the comfort of indoor environments (wang2024advancements, ).
Figure 2. Literature Review of Building Exteriors
[…]
Literature Review of Building Exteriors on Web of Science.
@@
In addition to fundamental building information and its spatial attributes, we have categorized the concept of ”smart building exterior design” into three primary areas that are most frequently discussed by scholars: ”human,” ”energy,” and ”green.” Incorporating these keywords into our search criteria, we have identified a total of 4,68346834,6834 , 683 relevant works, with the distribution being $544544$ for ”human,” 4,11541154,1154 , 115 for ”energy,” and $523523$ for ”green.”
@@
Given our focus on indicators that can be discerned through visual analysis, we have narrowed down the total literature on building exteriors to $5454$ relevant pieces. Within this subset, only $3939$ pieces specifically address the hot topic of human-centric, energy fuel and engineering environmental fields. We reviewed these works and we found that among these, the topic of glazing has garnered significant attention. Specific indicators such as glazing type, colour, and Window-Wall Ratio (WWR) are recognized as crucial factors influencing energy absorption (WWR, ; WWR-and-type, ). Outdoor air conditioners, as significant contributors to heat, and their placement methods are also central topics in discussions about energy conservation (aircon-position, ). The rooftop, as the uppermost exterior of the building, plays a pivotal role in capturing solar energy. Its design and materials can significantly influence the building’s energy efficiency by absorbing and utilizing solar radiation (energy-saving, ). From a human-centric perspective, building styles exhibit considerable variation, influenced by factors such as culture, climate, and history (ng2020perception, ). They are commonly recognized as falling into categories such as Historical, Somewhat Historical, No Significant Style, Somewhat Modern, and Modern (human-perception, ). Vertical greenery not only contributes to aesthetics but also has a positive impact on the thermal insulation of building facades (campiotti2022vertical, ). Meanwhile, the surrounding infrastructure, such as transport stations and parking lots, can significantly influence the design of building exteriors and is, in turn, affected by them (transport, ; parking-lots, ). Consequently, we will also consider these elements.
@@
Table 1. Building Exteriors Indicators
@@
Floor-to Floor height 3m, approximately 2m, unknown. inferences based on a prior knowledge
@@
Number of Floors 2, approximately 2, unknown. inferences based on a prior knowledge
@@
Street type residential street, local street, etc.
Neighbouring buildings 3 similar buildings, none, etc.
Greening conditions 3 trees, 3 grasslands, unknown, etc.
Street facilities 3 roads, 3 parking lots, unknown, etc.
Public transport 3 subway stations, 3 bus stops, none, unknown, etc.
Human perceptual ratings complex, original, ordered, pleasing, boring
@@
Exterior complexity complex, moderate, simple inferences based on a prior knowledge
@@
Streetscape perception scores Safer, wealthier, livelier, more beautiful, more depressing, more boring
3.2. Sampling of Buildings and Street View Imagery
@@
We begin by using the Nominatim API to fetch city names and their corresponding countries based on a query provided through the command line. Nominatim, an open-source geocoding service, outputs a list of cities and countries that match the query, allowing for accurate identification of the desired city and country(clemens2015geocoding, ). This makes it easier to directly locate the corresponding bounding box by simply inputting the city and country names, streamlining the data collection process.
For the building sampling, we employ the Overpass API to retrieve and save building data for the specified city and country. The Overpass API(olbricht2011overpass, ), which is used to query the OpenStreetMap (OSM) database, performs several tasks: it retrieves the bounding box coordinates for the given city and country, queries the OSM database to obtain building IDs and their coordinates within the bounding box, gathers additional details such as the address and height of each building, and saves the building data into a JSONL file, organized by building types.
@@
Finally, to match the buildings with their corresponding street view exteriors, we use the Google Street View API to download images for the locations specified in the JSONL file. This process involves creating a directory to store the images based on the name of the JSONL file, reading the latitude and longitude coordinates from the file, and downloading images using the Google Street View API for each specified location. The images are then saved in the created directory. To ensure accuracy, the building ID coordinates are centered within a 30-meter buffer, and the closest street view image is selected for each building, saved in a $600\times 300600 × 300$ format.
3.3. Annotation with Multimodal Large Language Model

### v2

@@ Abstract
Building exteriors Street-View imagery Multimodal large language model GeoAI Urban analytics.
11footnotetext: Corresponding authors and the first author(Zongrong Li) is a Master’s student.
1 Introduction
@@
Urban environments are increasingly becoming the focus of interdisciplinary research, where the integration of spatial and visual data is essential for understanding various aspects of urban life. Within this context, the study of building exteriors has gained significant attention due to its direct impact on critical areas such as energy efficiency[1], environmental sustainability[2], and human-centric urban design[3]. Specifically, analyzing building exteriors involves extracting detailed spatial and visual features, which are crucial for evaluating urban heat islands, energy consumption, and overall aesthetic coherence. Moreover, building exteriors serve not only as the physical interface between a structure and its environment but also play a crucial role in determining the overall energy performance and aesthetic appeal of urban spaces.
In this regard, Street View imagery has emerged as an effective tool for extracting detailed features of building exteriors, providing a unique perspective that traditional remote sensing methods often overlook. [4]. Unlike remote sensing imagery, which provides a top-down view of urban landscapes, street view imagery captures buildings at eye level, revealing intricate details of their spatial attributes, materials, and architectural styles[5]. This granular level of detail is vital for assessing the impact of building exteriors on factors such as urban heat islands, energy consumption, and the overall aesthetic coherence of neighborhoods. Therefore, integrating street view imagery with advanced feature extraction techniques is essential for a comprehensive analysis of building exteriors. While street view imagery offers these advantages, its potential remains underutilized in urban research, particularly when it comes to integrating this visual data with other spatial datasets to provide a comprehensive understanding of building exteriors and their broader urban implications. This highlights the growing need to harness street view imagery more effectively to address the multifaceted challenges of modern urban environments.
@@
Through a systematic search of the urban building exterior research, it is evident that there is a significant gap in the availability of comprehensive databases that integrate street view imagery with detailed building exterior information. Existing studies tend to focus on either the spatial characteristics of buildings—such as their geographic positioning and interactions with surrounding infrastructure—or the aesthetic and material aspects, including innovations in architectural design and material usage[6][7]. However, these perspectives are rarely combined to offer a full understanding of urban building exteriors. The lack of integrated databases that unify spatial data with visual and material information limits the ability to conduct in-depth analyses necessary for informed decision-making in urban planning, architectural design, and environmental policy. Additionally, while recent advancements have leveraged street view imagery for various urban analyses, they primarily focus on environmental factors or general urban features rather than systematically extracting and organizing building exterior information[8]. The integration of large language models with street view imagery for detailed urban analysis remains in its nascent stages, highlighting the need for a more integrated approach to understanding urban building exteriors.
@@
To address these knowledge deficits, we propose BuildingView, an innovative approach that integrates Street View imagery with multimodal large language models to create a comprehensive urban building exteriors database. By leveraging high-resolution visual data from Google Street View, combined with spatial and structural data obtained through the Overpass API and OpenStreetMap, this research sets a new standard for the integration of visual and spatial data in urban informatics, significantly advancing the field. Our key contributions include:
  1. 1.
    Systematic collection and organization of urban building exteriors indicators, providing a comprehensive understanding of critical factors such as energy efficiency, environmental sustainability, and human-centric design.
  2. 2.
    Development of BuildingView, a reusable tool that utilizes Street View imagery and multimodal large language models to construct detailed urban building exterior databases for any urban area globally.
  3. 3.
    Construction of extensive urban building exterior databases for New York, Amsterdam, and Singapore, enabling more informed decision-making in urban planning, architectural design, and environmental policy.
To outline the structure of this paper, the remainder is organized as follows. Section 2 provides a review of relevant literature on building exteriors, discussing both traditional and emerging research in this domain. Section 3 introduces our methodological framework, including the systematic literature review, the sampling approach using Overpass and Google Street View APIs, as well as the annotation process with large language models. Section 4 presents experimental details and results, followed by a comprehensive analysis of findings. Section 5 discusses the implications and limitations of our study. Finally, section 6 concludes the paper and offers perspectives for future research.
2 Related Work
The current section surveys prior research on building exteriors, discussing the indicators commonly examined in the context of urban informatics, architecture, and sustainability. We divide the related work into two major subsections. Subsection 2.1 reviews various building exterior indicators—ranging from structural attributes to energy and environmental aspects—while Subsection 2.2 focuses on how Street View Imagery (SVI) and large language models (LLMs) have been leveraged in urban building analysis. This structured overview helps in identifying knowledge gaps that our proposed approach aims to address.
2.1 Indicators of Building Exteriors Indicators
@@
In the realm of urban informatics and geoscience, the development of spatiotemporal databases pertaining to buildings has garnered significant attention. Researchers within the disciplines of urban planning and geography typically concentrate on the spatial characteristics of buildings, such as their geographic positioning and the dynamics of their interaction with adjacent infrastructure. For instance, the architectural and urban planning processes must meticulously consider the spatial interplay between buildings and their adjacent parking facilities to optimize energy efficiency [1].In contrast, architects and designers are more inclined to examine the intrinsic elements of buildings, including materials and aesthetic design. Recently, the architectural style and the materials used for the exterior of buildings are continually subject to innovation and enhancement [6, 7].
@@
The academic discourse has increasingly pivoted towards the examination of the connections between buildings and other indicators, with a particular emphasis on their impact on energy consumption and environmental sustainability. Especially, researchers are actively exploring ways to harness more energy from natural sources through building design, thereby reducing the energy costs associated with buildings. Additionally, they are working to mitigate the thermal impacts caused by the urban heat island effect, which is exacerbated by the heat generated within urban environments [9]. Concurrently, there is a burgeoning interest in the human-centric aspects of building design, particularly how it fosters engagement and interaction with the public from a participatory perspective [3].
@@
While numerous open-source databases exist that cater to specific domains of building information, they are often limited to a single aspect of building data. The foundational building attribute knowledge within these databases is typically derived from field surveys or footprint data extracted from remote sensing imagery [10]. To date, no open-source database has been developed that integrates street view imagery to provide a comprehensive, visually-driven repository encompassing the spatial. Thus, we want to build a spatial database with fundamental, and external surface properties of buildings, with a focus on their relevance to energy, environmental, and human-centric dimensions.
2.2 SVI and LLMs in Urban Building Analysis
@@
The advent of street view imagery (SVI) has revolutionized the way urban environments are analyzed, offering a rich and dynamic geospatial data source that rivals traditional remote sensing methods such as satellite imagery [5]. The comprehensive visual data captured by SVI allows for detailed analysis and understanding of urban environments at the street level, providing critical insights into various urban characteristics and phenomena.
@@
Previous research has demonstrated the utility of SVI in urban studies by integrating it with other urban data sources such as street networks, building information, demographic and socioeconomic data, and survey responses. For instance, SVI has been utilized to audit neighborhood environments [8], and analyze the effects of green view indices on pedestrian activity[11]. Additionally, studies have explored the use of SVI and machine learning to measure the built environment’s impact on crime [12], as well as investigating the relationship between neighborhood walkability factors and walking behaviors using big data approaches [13]. Furthermore, SVI has been applied at a global scale, as seen in the creation of extensive datasets that cover millions of street-level images across hundreds of cities, facilitating large-scale urban science and analytics[14].
@@ 3 Method
In this study, we introduce BuildingView, a novel framework for constructing urban building exterior databases using Street View imagery and multimodal large language models. The research framework is divided into three primary steps, as illustrated in Fig 1, subsection 3.1: a systematic literature review of building exterior indicators; subsection 3.2: sampling of buildings and street view imagery using the Overpass and Google Street View APIs to gather data points; and subsection 3.3: annotation of street view images using the ChatGPT-4O API to create the Urban Building Exteriors Database. Further details on each of these steps are provided below.
Figure 1: BuildingView Workflow: (A) Indicator Review; (B) Sampling; (C) Annotation with ChatGPT-4.0.
3.1 Systematic Literature Review of Building Exteriors Indicators
@@
Global building morphology indicators database has concluded many indicators for urban data science [15]. However, a database of buildings directly covering energy, green, and human-centred buildings has not yet been created. Our goal is to establish a building database that encompasses the majority of indicators of interest to researchers. This database will include comprehensive spatial information on buildings and knowledge extracted from street view images through advanced vision analytics and vision-based reasoning techniques.
@@
For that, we have done a systematic literature review on building exteriors Fig 2. The initial step is to identify emerging topics concerning building exteriors over the past five years. To achieve this, we conduct a keyword search on "building envelope," "exterior of building," and "building exterior" within the Web of Science database. Specifically, this search was performed on November 15, 2024, and restricted to publications from 2019 to 2024, yielding 6,28562856,2856 , 285 relevant scholarly works. There is a growing discussion about the intelligent design of external surfaces of buildings. The integration of innovative materials and intelligent technologies has the potential to significantly enhance the comfort of indoor environments [16].
Figure 2: Literature Review of Building Exteriors
@@
In addition to fundamental building information and its spatial attributes, we have categorized the concept of "smart building exterior design" into three primary areas that are most frequently discussed by scholars: "human," "energy," and "green." Incorporating these keywords into our search criteria, we have identified a total of 4,68346834,6834 , 683 relevant works, with the distribution being $544544$ for "human," 4,11541154,1154 , 115 for "energy," and $523523$ for "green."
@@
Given our focus on indicators that can be discerned through visual analysis, we have narrowed down the total literature on building exteriors to $5454$ relevant pieces. Within this subset, only $3939$ pieces specifically address the hot topic of human-centric, energy fuel and engineering environmental fields. We reviewed these works and we found that among these, the topic of glazing has garnered significant attention. Specific indicators such as glazing type, colour, and Window-Wall Ratio (WWR) are recognized as crucial factors influencing energy absorption [17, 18]. Outdoor air conditioners, as significant contributors to heat, and their placement methods are also central topics in discussions about energy conservation [19]. The rooftop, as the uppermost exterior of the building, plays a pivotal role in capturing solar energy. Its design and materials can significantly influence the building’s energy efficiency by absorbing and utilizing solar radiation [20]. From a human-centric perspective, building styles exhibit considerable variation, influenced by factors such as culture, climate, and history [21]. They are commonly recognized as falling into categories such as Historical, Somewhat Historical, No Significant Style, Somewhat Modern, and Modern [3]. Vertical greenery not only contributes to aesthetics but also has a positive impact on the thermal insulation of building facades [22]. Meanwhile, the surrounding infrastructure, such as transport stations and parking lots, can significantly influence the design of building exteriors and is, in turn, affected by them [23, 1]. Consequently, we will also consider these elements.
@@
Table 1: Building Exteriors Indicators
@@
Floor-to-Floor height 3m, approximately 2m, unknown. inferences based on prior knowledge
@@
Number of Floors 2, approximately 2, unknown. inferences based on prior knowledge
@@
Street type residential street, local street, etc. -
Neighbouring buildings 3 similar buildings, none, etc. -
Greening conditions 3 trees, 3 grasslands, unknown, etc. -
Street facilities 3 roads, 3 parking lots, unknown, etc. -
Public transport 3 subway stations, 3 bus stops, none, unknown, etc. -
Human perceptual ratings complex, original, ordered, pleasing, boring -
@@
Exterior complexity complex, moderate, simple inferences based on prior knowledge
@@
Streetscape perception scores Safer, wealthier, livelier, more beautiful, more depressing, more boring -
3.2 Sampling of Buildings and Street View Imagery
@@
We begin by using the Nominatim API to fetch city names and their corresponding countries based on a query provided through the command line. Nominatim, an open-source geocoding service, outputs a list of cities and countries that match the query, allowing for accurate identification of the desired city and country[24]. This makes it easier to directly locate the corresponding bounding box by simply inputting the city and country names, streamlining the data collection process.
For the building sampling, we employ the Overpass API to retrieve and store building data for the specified city and country. The Overpass API [25], which queries the OpenStreetMap (OSM) database, returns building footprints, latitude/longitude coordinates, and occasional attributes (e.g., address, height). While this information is valuable for establishing each building’s spatial context, OSM data generally lacks detailed façade descriptors—such as window-to-wall ratio or materials—that are essential for our analysis. Consequently, we integrate street-view imagery to capture more fine-grained exterior attributes. The Overpass-sourced building records are saved in JSONL format by building types, providing a structured foundation for subsequent merging with street-view data.
@@
Finally, to match the buildings with their corresponding street view exteriors, we use the Google Street View API to download images for the locations specified in the JSONL file. This process involves creating a directory to store the images based on the name of the JSONL file, reading the latitude and longitude coordinates from the file, and downloading images using the Google Street View API for each specified location. To ensure that each downloaded street-view image captures the correct building exterior (rather than facing the opposite side of the street), we adopt the following strategy to determine the camera orientation (heading) when querying the Google Street View API. First, we compute the bearing from the street-view panorama’s GPS coordinate to the building’s centroid. Specifically, we retrieve the panorama’s latitude and longitude returned by the Street View API (the closest panorama to our target building), and then calculate the azimuth angle toward the building’s coordinates. This angle is subsequently used as the heading parameter in our Street View API request, ensuring that the image is oriented to face the building.
Additionally, if no panorama is found within $3030$ meters of the building’s coordinate, we expand the search radius in incremental steps (e.g., up to $5050$ meters), ensuring a higher likelihood of acquiring at least one valid image in dense urban areas. Through this bearing-based approach, we minimize errors where the camera might otherwise capture the opposite side of the street or irrelevant surroundings. The final images, saved at $600\times 300600 × 300$ resolution, are thus more likely to depict the intended building facade.
3.3 Annotation with Multimodal Large Language Model